### v1

@@ 1 Introduction
An important subtlety arises when realizing higher-form symmetries in lattice models with tensor product Hilbert spaces. In such settings, higher-form symmetry generators are not strictly topological operators at the microscopic level. Instead, they are emergent and become topological operators only in the presence of an energetically enforced Gauss law at low energies. This emergent nature complicates the definition and characterization of ’t Hooft anomalies. An ’t Hooft anomaly refers to an obstruction to gauging a global symmetry, and its presence implies that the system cannot be realized in a short-range entangled (SRE) phase [2]. For instance, the Lieb-Schultz-Mattis theorem [3, 4, 5] and its generalizations [6, 7, 8, 9, 10, 11, 12, 13, 14] encode non-trivial constraints on the low-energy physics of lattice systems that originate from mixed anomalies between spatial and internal symmetries. Anomalies also put strong constraints on deconfinement in gauge theories [15, 16, 17] and imply the nontrivial edge states of symmetry protected topological phases [18, 19, 20, 21]. ’t Hooft anomalies of higher-form symmetries have recently been used to constrain the entanglement structure of both pure and mixed states, providing refined dynamical constraints on the system [22, 23, 24, 25, 26].
@@ 2.2 Else-Nayak type index
    where $p\in R$ denotes the plaquettes along the boundary $\partial R$, and we choose a restriction $W_{p;R}$ of the Gauss law operator. $p\in\text{Int}(R)$ denotes the plaquettes inside the region, where $W_{p}$’s are not truncated. A pair of operators $W_{p;R}$ at neighboring plaquettes along the boundary $\partial R$ no longer commute with each other. Therefore, we are fixing an ordering of operators to define the product $\prod_{p\in\partial R}$, such that $U_{R}(\epsilon)$ is still interpreted as a finite-depth circuit. Later in Sec. 2.2.2, we will see that the anomaly index is independent of possible ambiguities to define $U_{R}(\epsilon)$, i.e., the ordering in the product and choices of truncations to define $W_{p;R}$.
@@ 3 Higher-form symmetry and anomaly
[…]
@@ 3.1.1 Warm-up: 1-form symmetry
This allows us to perform “dimensional reduction” of the symmetry operator. Let us take a $d$-dimensional disk $D^{d}$ in the triangulation $\Lambda$, then we define a restriction of the symmetry operator within $R$ by
@@ 5 Conclusions
  • For 0-form symmetry in (1+1)D, is known that being free of anomalies is equivalent to being onsite-able [44]. Meanwhile, higher-form symmetry can be anomalous while being onsite as discussed in Sec. 2.2.3. It would be interesting to understand the criteria for onsiteability of generic higher-form symmetries.
@@ Appendix A Review: Else-Nayak index for 0-form symmetry in (1+1)D
$\Gamma_{\partial I}(g,h)\,\Gamma_{\partial I}(gh,k)\;=\;\Gamma_{\partial I}(g,hk)\,(^{g}\Gamma_{\partial I}(h,k))~.$ (82)
@@
$\Gamma_{l}(g,h)\,\Gamma_{l}(gh,k)\;=\;\omega(g,h,k)\,\Gamma_{l}(g,hk)\,(^{g}\Gamma_{l}(h,k))~,$ (83)
@@
$\omega(h,k,\ell)\,\omega(g,hk,\ell)^{-1}\,\omega(g,h,k\ell)\,\omega(g,h,k)^{-1}\;=\;1,$ (84)
so that $\omega\in Z^{3}(G,U(1))$. Different choices of $\Gamma_{l}$ by a phase $\chi(g,h)$ shifts $\omega$ by a coboundary, and the anomaly is uniquely characterized by the cohomology class

### v2

@@ 1 Introduction
An important subtlety arises when realizing higher-form symmetries in lattice models with tensor product Hilbert spaces. In such settings, higher-form symmetry generators are not strictly topological operators at the microscopic level. Instead, they are emergent and become topological operators only in the presence of an energetically enforced Gauss law at low energies. This emergent nature complicates the definition and characterization of ’t Hooft anomalies. An ’t Hooft anomaly refers to an obstruction to gauging a global symmetry, and its presence implies that the system cannot be realized in a short-range entangled (SRE) phase [2]. For instance, the Lieb-Schultz-Mattis theorem [3, 4, 5] and its generalizations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] encode non-trivial constraints on the low-energy physics of lattice systems that originate from mixed anomalies between spatial and internal symmetries. Anomalies also put strong constraints on deconfinement in gauge theories [17, 18, 19] and imply the nontrivial edge states of symmetry protected topological phases [20, 21, 22, 23]. ’t Hooft anomalies of higher-form symmetries have recently been used to constrain the entanglement structure of both pure and mixed states, providing refined dynamical constraints on the system [24, 25, 26, 27, 28].
@@ 2.2 Else-Nayak type index
    where $p\in R$ denotes the plaquettes along the boundary $\partial R$, and we choose a restriction $W_{p;R}$ of the Gauss law operator. $p\in\text{Int}(R)$ denotes the plaquettes inside the region, where $W_{p}$’s are not truncated. A pair of operators $W_{p;R}$ at neighboring plaquettes along the boundary $\partial R$ no longer commute with each other. Therefore, we are fixing an ordering of operators to define the product $\prod_{p\in\partial R}$. [^2] Due to mesoscopic nature of the lattice $\hat{\Lambda}$, $U_{R}(\epsilon)$ is a finite-depth circuit with any choice of ordering in the product. Later in Sec. 2.2.2, we will see that the anomaly index is independent of possible ambiguities to define $U_{R}(\epsilon)$, i.e., the ordering in the product and choices of truncations to define $W_{p;R}$.
@@ 2.2.5 Graphical representation of Else-Nayak type index
2.3 Example: ${\mathbb{Z}}_{2}$ toric code in (2+1)D
As an explicit example, let us compute the anomaly index $\omega_{4}]\in H^{4}(B^{2}G,U(1))$ for the ${\mathbb{Z}}_{2}$ 1-form symmetry of the (2+1)D ${\mathbb{Z}}_{2}$ toric code. This ${\mathbb{Z}}_{2}$ 1-form symmetry is onsite, given by a product of Pauli $X,Z$ operators along a string.
Let us label the vertices of $\Lambda$ at $\partial R$ by numbers $\ldots,j-1,j,j+1,\ldots$, so that each truncated Gauss law operator $W_{p;R}$ is simply denoted by $W_{j}$, where a plaquette $p$ corresponds to a vertex $j$ of $\Lambda$. See Fig. 4 for an illustration. Since the 1-form symmetry is onsite, each truncated Gauss law operator $W_{j}$ again satisfies the ${\mathbb{Z}}_{2}$ algebra,
$(W_{j})^{2}=1~.$ (32)
Also, reflecting that the string operator is a fermion, $W_{j}$ follows the commutation relation
$W_{j}W_{j+1}=-W_{j+1}W_{j}~.$ (33)
Let us now compute the $H^{4}$ index of this ${\mathbb{Z}}_{2}$ 1-form symmetry. The truncated symmetry operator $W_{R}(\epsilon)$ has the form of
$U_{R}(\epsilon)=\left(...W_{j-1}^{\epsilon(j-1)}W_{j}^{\epsilon(j)}W_{j+1}^{\epsilon(j+1)}...\right)\times\prod_{p\in\text{Int}(R)}(W_{p})^{\epsilon(p)}~.$ (34)
Using (32), (33), the reduced operator $\Omega(\epsilon_{01},\epsilon_{12},g_{012})=U_{R}(\epsilon_{01})U_{R}(\epsilon_{12})U_{R}(\epsilon_{01}+\epsilon_{12}-g_{012})^{-1}$ is expressed as
$\Omega(\epsilon_{01},\epsilon_{12},g_{012})=(-1)^{\sum_{j}(\epsilon_{01}(j)+\epsilon_{12}(j))g_{012}+\epsilon_{01}(j+1)\epsilon_{12}(j)}\times U_{R}(g_{012})~.$ (35)
The contribution from $U_{R}(g_{012})$ provides only an overall factor in $F_{e}$ that is independent of the 0-cochains $\epsilon$ and hence does not affect the $H^{4}$ index. We therefore focus on the remaining phase factor
$O_{e}(\epsilon_{01},\epsilon_{12},g_{012}):=(-1)^{\sum_{j}\bigl(\epsilon_{01}(j)+\epsilon_{12}(j)\bigr)g_{012}\;+\;\epsilon_{01}(j+1)\epsilon_{12}(j)}~,$ (36)
which depends only on the local values of $\epsilon_{01}$ and $\epsilon_{12}$ near the edge $e=\langle j,j+1\rangle\subset\partial R$. From this local functional $O_{e}$, we obtain
$\begin{split}F_{e}(\epsilon_{01},\epsilon_{12},\epsilon_{23},\{g\})&=\frac{1}{2}\left((\epsilon_{01}(j)+\epsilon_{12}(j))g_{012}+\epsilon_{01}(j+1)\epsilon_{12}(j)+(\epsilon_{02}(j)+\epsilon_{23}(j))g_{023}+\epsilon_{02}(j+1)\epsilon_{23}(j)\right)\\ &-\frac{1}{2}\left((\epsilon_{12}(j)+\epsilon_{23}(j))g_{123}+\epsilon_{12}(j+1)\epsilon_{23}(j)+(\epsilon_{01}(j)+\epsilon_{13}(j))g_{013}+\epsilon_{01}(j+1)\epsilon_{13}(j)\right)~.\end{split}$ (37)
By a repeated use of $\epsilon_{ik}=\epsilon_{ij}+\epsilon_{jk}-g_{ijk}$ and $g_{013}+g_{123}=g_{012}+g_{023}$ mod 2, the above expression of $F_{e}$ is rewritten as
$F_{e}(\epsilon_{01},\epsilon_{12},\epsilon_{23},\{g\})=\frac{1}{2}(\epsilon_{01}(j)~g_{123}+\epsilon_{01}(j+1)~g_{123})+\phi(\{g\})\quad\mod 1~,$ (38)
where $\phi(g)$ is a functional that solely depends on $\{g\}$. We again ignore the factor $\phi(\{g\})$ independent of 0-cochains $\epsilon$ since it does not affect the computation of the $H^{4}$ index. Then $A_{j}$ with $F=dA$ is given by
$A_{j}=\frac{1}{2}\epsilon_{01}(j)~g_{123}~.$ (39)
Finally, the index $\omega_{4}$ is given by
$\begin{split}\omega_{l}(\epsilon_{01},\epsilon_{12},\epsilon_{23},\epsilon_{34},\{g\})&=A_{j}(\epsilon_{01},\epsilon_{12},\epsilon_{23})+A_{j}(\epsilon_{01},\epsilon_{13},\epsilon_{34})+A_{j}(\epsilon_{12},\epsilon_{23},\epsilon_{34})-A_{j}(\epsilon_{02},\epsilon_{23},\epsilon_{34})-A_{j}(\epsilon_{01},\epsilon_{12},\epsilon_{24})~\\ &=\frac{1}{2}\epsilon_{01}~(g_{123}+g_{134}-g_{124})+\frac{1}{2}\epsilon_{12}~g_{234}-\frac{1}{2}\epsilon_{02}~g_{234}\\ &=\frac{1}{2}g_{012}~g_{234}~.\end{split}$ (40)
Therefore we get
$\omega_{4}(\{g\})=\frac{1}{2}g_{012}~g_{234}~,$ (41)
which can be written as $\frac{1}{2}g\cup g$ evaluated at a 4-simplex $\langle 01234\rangle$, by regarding $g$ as a 2-cocycle $g\in Z^{2}(B^{2}G,{\mathbb{Z}}_{2})$. Here, $\cup$ is a cup product of cochains on the triangulation $\Lambda$, defined as
$a_{k}\cup a_{l}(0,\dots,k+l)=a_{k}(0,\dots,k)a_{l}(k,\dots,l)~,$ (42)
with $k,l$-cochains $a_{k},a_{l}$. The cup product generally induces a map between cohomology classes, $\cup:H^{k}\times H^{l}\to H^{k+l}$. Explicitly, $g$ is a $\mathbb{Z}_{2}$-valued $2$-cocycle taking values 0 or $1$; consequently, $\omega_{4}$ takes values 0 or $\tfrac{1}{2}$ in Eq. (41), where we embed ${\mathbb{Z}}_{2}$ in $U(1)=\mathbb{R}/\mathbb{Z}$. Therefore $\omega$ defines an element of $Z^{4}(B^{2}G,U(1))$, producing the desired (3+1)D response action for the anomaly [36, 37, 38].
Figure 4: (a): The truncated symmetry operator $W_{j}$ of the ${\mathbb{Z}}_{2}$ toric code. A single qubit is located at each edge of a square lattice, and the operator is a product of Pauli $X$ and $Z$ operators. (b): At the boundary of the region $R$, there is an array of truncated operators $\ldots,W_{j-1},W_{j},W_{j+1},\ldots$ The neighboring operators anti-commute with each other due to fermionic statistics.
@@ 3 Higher-form symmetry and anomaly
with a (p2)(p-2)-cochain $\eta\in C^{p-2}(\Lambda,G)$. The property (45) is not well-defined for $1$-form symmetry, $p=1$. In the case of $p=1$, we instead assume the property
@@ 3.1.1 Warm-up: 1-form symmetry
This allows us to perform “dimensional reduction” of the symmetry operator. Let us take a $d$-dimensional disk $D^{d}$ in the triangulation $\Lambda$, then we define a restriction of the symmetry operator within $D^{d}$ by
@@ 5 Conclusions
  • For 0-form symmetry in (1+1)D, is known that being free of anomalies is equivalent to being onsite-able [46]. Meanwhile, higher-form symmetry can be anomalous while being onsite as discussed in Sec. 2.2.3. It would be interesting to understand the criteria for onsiteability of generic higher-form symmetries  [^5] Recently, some of the present authors have identified an index that characterizes the obstruction to onsiteability of 1-form symmetries in (3+1)D, taking values in $H^{3}(B^{2}G,\mathbb{Q}_{+})$ [47]..
@@ Appendix A Review: Else-Nayak index for 0-form symmetry in (1+1)D
$\Gamma_{\partial I}(g,h)\Gamma_{\partial I}(gh,k)\;=\;(^{g}\Gamma_{\partial I}(h,k))\Gamma_{\partial I}(g,hk)~.$ (92)
@@
$\Gamma_{l}(g,h)\,\Gamma_{l}(gh,k)\;=\;\omega(g,h,k)\,(^{g}\Gamma_{l}(h,k))\,\Gamma_{l}(g,hk)~,$ (93)
@@
$\omega(g,h,k)\omega(g,hk,l)\omega(h,k,l)\left(\omega(gh,k,l)\omega(g,h,kl)\right)^{-1}=1~,$ (94)
so that $\omega\in Z^{3}(BG,U(1))$. Different choices of $\Gamma_{l}$ by a phase $\chi(g,h)$ shifts $\omega$ by a coboundary, and the anomaly is uniquely characterized by the cohomology class